OpenStudy (anonymous):

the graph of a quadratic function (a parabola) has x-intercepts -1 and 3 and a range consisting of all numbers less than or equal to 4. determine an expression for the function

OpenStudy (dumbcow):

factored form y = a(x+1)(x-3) max = 4 find x_coord of vertex Because of symmetry, vertex is exactly in the middle of x=-1 and x=3 (-1+3)/2 = 2/2 = 1 vertex: (1,4) plug in these x,y values to solve for a 4 = a(1+1)(1-3) 4=a(2)(-2) 4 = -4a -1 = a Final equation: y = -(x+1)(x-3)

OpenStudy (anonymous):

thank you!

OpenStudy (dumbcow):

no problem hope that made sense